OpenStudy (anonymous):

Using the ratio of perfect squares method, what is the square root of 53 rounded to the nearest hundredth?

OpenStudy (anonymous):

and calculators would be wrong in this, apparently

OpenStudy (anonymous):

@AkashdeepDeb @iGreen

OpenStudy (jollypikachu300):

the anwser would be 53.00

OpenStudy (anonymous):

Can you show me how you got this answer?

OpenStudy (jollypikachu300):

10 times 53=53.00

OpenStudy (tkhunny):

Why would you want to know how to get that answer? 53 is NOT the square root of 53.

OpenStudy (anonymous):

Oooooh....I feel dumb now LOL

OpenStudy (jollypikachu300):

yes

OpenStudy (tkhunny):

This is called "Linear Interpolation" 7^2 = 49 8^2 = 64 (53-49)/(64-49) = (SR - 7)/(8-7) 4/15 = (SR-7) SR = 7 + 4/15 = 7.266666 7.266666^2 = 52.804434 <== So we're still a little short. 7.266666^2 = 52.804434 8^2 = 64 (53-52.804434)/(64-52.804434) = (SR - 7.266666)/(8-7.266666) 0.195565/11.195566 = (SR-7.266666) SR = 7.266666 + 0.017468076 = 7.284134 7.284134^2 = 53.05861 <== We overshot a little.

OpenStudy (tkhunny):

Do you have an example of this "Ratio of Perfect Squares" method? Maybe an example from your text book?